OpenStudy (anonymous):

Solve x + 3y + z = 1 System of Equations x, y, z 2x – y – z = 6 5x + y + z = 1

OpenStudy (anonymous):

Answers: (1, 6, 1) (-1, -2, 6) (1, 2, -6) (2, 1, -4)

OpenStudy (anonymous):

I meant to ask you for help on this one too

OpenStudy (anonymous):

This is basically another question that's exactly the same format as the one I solved before. Please look at what I did in the previous question to answer this one yourself.

OpenStudy (anonymous):

Start by getting two equations with two variables. To do that, choose a variable to eliminate. HINT: Choose z, because you don't need to multiply to get rid of it in all three equations.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what is the next step?

OpenStudy (anonymous):

Show me your two equations that have only two variables in both equations first.

OpenStudy (anonymous):

2x - y -z = 6 +5x = y + z= 1

OpenStudy (anonymous):

*+

OpenStudy (anonymous):

Do those two equations have two variables or three? Didn't I tell you to get two equations that have only TWO VARIBALES?

OpenStudy (anonymous):

VARIABLES*

OpenStudy (anonymous):

Example?

OpenStudy (anonymous):

When someone like #iforget asks you if you understand how to solve the previous question, you answer honestly. If you don't understand then please say no and not just yes and go to the next question. If you were to do a written exam, teachers don't mark your test to see whether you get the correct answer or not. Multiple choice is practice for you to know how to get the answer. Not just randomly guess at the answer.

OpenStudy (anonymous):

And an example was shown to you in the previous question...

OpenStudy (anonymous):

If you read what I did in the previous question, you would know an example that had two variables.

OpenStudy (anonymous):

I meant to save it but I closed the question

OpenStudy (anonymous):

Then go to the closed section and read what I wrote.

OpenStudy (anonymous):

Ok I see

OpenStudy (anonymous):

There's two sections: Open Questions and Closed Questions.

OpenStudy (anonymous):

Yea I barely use this. I got it thanks

OpenStudy (anonymous):

No worries.